OpenStudy (yessgeorgie8):

15=5^x+3 -7. Solve for x, round to nearest thousandths PLEASE AND THANK YOU

OpenStudy (anonymous):

i take it it is \[15=5^{x+3}-7\] right ?

OpenStudy (yessgeorgie8):

yes @satellite73

OpenStudy (anonymous):

step one add 7 to both sides, what do you get ?

OpenStudy (yessgeorgie8):

22=5^x+3

OpenStudy (yessgeorgie8):

@satellite73

OpenStudy (anonymous):

yes, then use the change of base formula do you know it?

OpenStudy (yessgeorgie8):

x+3=log5 22 ?? @satellite73

OpenStudy (anonymous):

yes but that won't help you get a number, that is equivalent logarithmic form

OpenStudy (anonymous):

it says "nearest thousandths" so they want a decimal out of you

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] is the change of base formula, something you can put in a calculator to get an answer

OpenStudy (anonymous):

you got \[5^{x+3}=22\] so \[x+3=\frac{\ln(22)}{\ln(5)}\]

OpenStudy (yessgeorgie8):

i got

OpenStudy (yessgeorgie8):

x+3=1.920

OpenStudy (anonymous):

i didn't do it but i am sure you can

OpenStudy (yessgeorgie8):

thank you thank u i got it ! i have one more question please??

OpenStudy (anonymous):

sure

OpenStudy (yessgeorgie8):

9x^2-6x+5=0

OpenStudy (yessgeorgie8):

describe the solutiona for this quadratic equation

OpenStudy (yessgeorgie8):

solutions

OpenStudy (anonymous):

describe as in solve or describe as in "two real, one real, two complex"?

OpenStudy (yessgeorgie8):

describe in as if it has, 2 rational roots, 2 complex roots, 1 rational double root, or 2 irrational roots

OpenStudy (anonymous):

probably the second one compute the discriminant \[b^2-4ac\] for \(a=9,b=-6,c=5\)

OpenStudy (yessgeorgie8):

got 216

OpenStudy (anonymous):

it if is a perfect square you get two rational ones if it is positive but not a perfect square they are irrational, it if is negative they are complex, and if it is 0 there is one rational double root

OpenStudy (yessgeorgie8):

so confused, so would it be 2 complex roots??

OpenStudy (anonymous):

yes, but the discriminant is not what you wrote

OpenStudy (anonymous):

\[(-6)^2-4\times 9\times 5=?\]

OpenStudy (yessgeorgie8):

-144?

OpenStudy (anonymous):

yes negative, so two complex roots

OpenStudy (yessgeorgie8):

thank you!